How much heat would be required to convert 173.3 g of ice to water at 0°C? (AHfus = 6.01 kJ/mol for water) 7 8 +/- x 100
Answer #1

Molar mass of H2O,

MM = 2*MM(H) + 1*MM(O)

= 2*1.008 + 1*16.0

= 18.016 g/mol

mass(H2O)= 173.3 g

use:

number of mol of H2O,

n = mass of H2O/molar mass of H2O

=(1.733*10^2 g)/(18.02 g/mol)

= 9.619 mol

Now use:

Q = number of mol * Hfus

= 9.619 mol * 6.01 KJ/mol

= 57.8 KJ

Answer: 57.8 KJ